OpenStudy (anonymous):

Choose the BLANK step in the given solution to the inequality -2x - 14 > -4 + 3x. -2x - 14 > -4 + 3x -5x - 14 > -4 BLANK x < -2 Answer -5x < 10 -5x > -28 x < 2 -5x > 10

OpenStudy (karatechopper):

last one

OpenStudy (anonymous):

Choose the BLANK step in the given solution to the inequality 6x - 2<= 8x + 14. 6x - 2<= 8x + 14 -2x - 2 <= 14 -2x<= 16 BLANK x>= 8 x >= -8 x <= -8 x <= 8

OpenStudy (mr.math):

karate was right about your first problem.

OpenStudy (karatechopper):

hehehe

OpenStudy (anonymous):

Yes she was thanks guys

OpenStudy (mr.math):

Sorry, second one :D

OpenStudy (karatechopper):

welcome